# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 4.2 Flux functionals on discrete surfaces
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=\sum_{f\in F(\Sigma_{h,i})}\int_{f}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=\int_{\partial\Sigma_{h,j}}{\bm{\tau}}\cdot{\bm{a}}_{h},\qquad{\bm{a}}_{h}\in\tilde{V}^{1}_{h},$
where the last equality uses the fact that integrals on interior curves, i.e., curves of the type $\partial f\cap\partial f^{\prime}$ with $f\neq f^{\prime}$, cancel because of their opposite orientation and the tangential continuity of the fields in $H(\operatorname{{\bf curl}};\Omega)$. As a consequence, we obtain
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{a}}_{h}=0$ (57)
for all ${\bm{a}}_{h}\in\tilde{V}^{1}_{0,h}$, and
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{a}}=\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot{\bm{a}}$ (58)
for all ${\bm{a}}\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{1}$.
@@ 4.3 Discrete harmonic spaces
A discretization of the harmonic spaces (5) is classically provided by the harmonic spaces associated with the discrete complex [4]. Since the primal and dual complexes have the same harmonic spaces at the continuous level, one can use the dual discrete complex (54): this leads to the discrete vector-valued spaces
@@
where the discrete kernel spaces are defined as usual,
@@
$\mathfrak{H}^{1}_{h}:=\tilde{\mathfrak{H}}^{2}_{h,0},\qquad\mathfrak{H}^{2}_{h}:=\tilde{\mathfrak{H}}^{1}_{h,0}.$ (61)
@@
This property holds for Whitney, Nedelec and general FEEC spaces [4, Sec. 5.5]. Finally, we can verify that discrete potentials may a priori be used for the discrete harmonic fields.
@@
Let ${\bm{v}}_{h}\in\tilde{\mathfrak{H}}^{1}_{h,0}$. We have ${\bm{v}}_{h}\in H_{0}(\operatorname{{\bf curl}}0;\Omega)$ by construction and the argument used to show (21) yields ${\bm{v}}_{h}=\operatorname{{\bf grad}}\phi$ for some $\phi\in H^{1}(\Omega)$. Since ${\bm{v}}_{h}$ is bounded as a piecewise continuous function, we observe that $\phi$ is also (Lipschitz) continuous, in particular we may apply the commuting projecion operators (55). This yields
@@
$\bar{\bm{A}}({\bm{x}})=\Big{(}0,\int_{-R}^{x_{1}}\bar{w}_{h,3}(x^{\prime},x_{2},x_{3})\,{\rm d}x^{\prime},-\int_{-R}^{x_{1}}\bar{w}_{h,2}(x^{\prime},x_{2},x_{3})\,{\rm d}x^{\prime}\Big{)},$
where $\bar{\bm{w}}_{h}$ is defined as the extension of ${\bm{w}}_{h}$ by 0 on $C_{R}\setminus\Omega$. From ${\bm{w}}_{h}\in C^{0}(\mathcal{M}_{h})$ we infer that ${\bm{A}}\in C^{0}(\mathcal{M}_{h})\cap H(\operatorname{{\bf curl}};\Omega)$, so that we may again apply the commuting projection operators (55). This yields
@@ 4.4 Discrete harmonic scalar potentials for $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$
The fields ${\bm{v}}_{h,i}:=\operatorname{{\bf grad}}\phi_{h,i}$, $i=1,\dots,\beta_{2}$, form a basis of $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$. Moreover, they satisfy
@@ 4.5 Discrete harmonic vector potentials for $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$
$\left\{\begin{aligned} \langle\sigma_{h,i},\tau\rangle+\langle\operatorname{{\bf grad}}\tau,{\bm{A}}_{h,i}^{0}\rangle&=0,\ &&\quad\forall\ \tau\in\tilde{V}^{0}_{h,0},\\ \langle\operatorname{{\bf grad}}\sigma_{h,i},{\bm{v}}\rangle+\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{0},\operatorname{{\bf curl}}{\bm{v}}\rangle+\langle{\bm{p}}_{h,i},{\bm{v}}\rangle&=-\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{b},\operatorname{{\bf curl}}{\bm{v}},,\rangle\ &&\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0},\\ \langle{\bm{A}}_{h,i}^{0},{\bm{q}}\rangle&=0,\ &&\quad\forall\ {\bm{q}}\in\tilde{\mathfrak{H}}^{1}_{h,0}.\end{aligned}\right$. (74)
@@
Observing that the flux functionals $\Phi_{i}$ from (34) are well-defined on $\tilde{\mathfrak{H}}^{2}_{h,0}$ since the latter is a subspace of $\tilde{V}^{2}_{0}=H_{0}(\operatorname{div};\Omega)$, we then have the following result.
@@
The fields ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$, $i=1,\dots,\beta_{1}$, form a basis of $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$. Moreover, their fluxes on the surfaces introduced in Section 4.2 satisfy
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot{\bm{w}}_{h,i}=\delta_{i,j},\quad 1\leq i,j\leq\beta_{1}.$ (75)
@@
The proof is now a matter of simple verifications. By taking ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{h,i}$ in (74) we find that $\sigma_{h,i}=0$, and with ${\bm{v}}={\bm{p}}_{h,i}$ we obtain ${\bm{p}}_{h,i}=0$. The first and last equations from (74) yield then
@@
(where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (59)–(60)), while the second equation from (74) gives
$\langle\operatorname{{\bf curl}}({\bm{A}}^{0}_{h,i}+{\bm{A}}^{b}_{h,i}),\operatorname{{\bf curl}}{\bm{v}}\rangle=0,\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0}.$ (77)
In particular, ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$ satisfies $\operatorname{div}{\bm{w}}_{h,i}=0$ and ${\bm{w}}_{h,i}\in(\operatorname{{\bf curl}}\tilde{V}^{1}_{h,0})^{\perp}$. As for the boundary condition ${\bm{n}}\cdot{\bm{w}}_{h,i}=0$ on $\partial\Omega$, it is verified by writing
where we have used (73) and the commutation property (55) in the second equality. Here both term satisfies the boundary condition: the first one thanks to (38) and (53), and the second one as the curl of a field in $H_{0}(\operatorname{{\bf curl}};\Omega)$. Thus we have verified that the fields ${\bm{w}}_{h,i}$ all belong to $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$ indeed.
To evaluate the fluxes across the surfaces $\Sigma_{h,j}$, we use (57) and (58): this yields
$\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{h,i}=\int_{\Sigma_{h,j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}^{b}_{h,i}=\int_{\partial\Sigma_{h,j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\delta_{i,j}$
where in the last two steps we have used the fact that the curves $\partial\Sigma_{h,j}$ and $\partial\Sigma_{j}$ coincide on the support of ${\bm{A}}^{b}_{i}$, see (56), and the boundary conditions (38). This shows (75) and the linear independence of the fields ${\bm{w}}_{h,i}$. Their spanning properties follow from (62). ∎
@@ 4.6 Solving a simpler discrete curl-curl problem
where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}=\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}\oplus\tilde{\mathfrak{H}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (59)–(60). Since the relations (80) characterize a unique field in $\tilde{V}^{1}_{h,0}$, this will show the existence and uniqueness of a solution. The fact that it coincides with the solution of (74) will follow from the fact that the latter satisfies the same relations, as seen in the proof of Theorem 4.2.

### v2

@@ 4.1 Discretization of the homogeneous and extended dual complexes
$\int_{\Sigma_{i}}{\bm{n}}\cdot\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{a}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot\tilde{\Pi}^{1}{\bm{a}}=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{a}},\qquad{\bm{a}}\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1}.$ (69)
@@ 4.2 Discrete harmonic spaces
A discretization of the continuous harmonic spaces (5) is classically provided by the harmonic spaces of the discrete complex [6]. Since the primal and dual complexes have the same harmonic spaces at the continuous level, one may use the dual discrete complex (67): this leads to the discrete vector-valued spaces
@@
where the discrete kernels are defined as usual,
@@
$\mathfrak{H}^{1}_{h}:=\tilde{\mathfrak{H}}^{2}_{h,0},\qquad\mathfrak{H}^{2}_{h}:=\tilde{\mathfrak{H}}^{1}_{h,0}.$
Remark 4.1.
The choice made here of discretizing the tangent harmonic spaces $\mathfrak{H}^{1}_{h}$ with 2-forms is consistent with the use of (strong) vector potentials discretized as 1-forms. A different choice is made in [1], where the harmonic fields of $\mathfrak{H}^{1}$ are represented by discrete loop fields, which are discrete 1-forms with nonzero circulations along 1-cycles that correspond to our tunnel curves.
@@
This property holds for Whitney, Nedelec and general FEEC spaces [6, Sec. 5.5]. Finally, we can verify that discrete potentials may a priori be used for the discrete harmonic fields.
@@
Let ${\bm{v}}_{h}\in\tilde{\mathfrak{H}}^{1}_{h,0}$. We have ${\bm{v}}_{h}\in H_{0}(\operatorname{{\bf curl}}0;\Omega)$ by construction and the argument used to show (21) yields ${\bm{v}}_{h}=\operatorname{{\bf grad}}\phi$ for some $\phi\in H^{1}(\Omega)$. Since ${\bm{v}}_{h}$ is piecewise smooth (and bounded), we observe that $\phi$ is also piecewise smooth, in particular we may apply the commuting projecion operators (68). This yields
@@
$\bar{\bm{A}}({\bm{x}})=\Big(0,\int_{-R}^{x_{1}}\bar{w}_{h,3}(x^{\prime},x_{2},x_{3})\ \mathrm{d}x^{\prime},-\int_{-R}^{x_{1}}\bar{w}_{h,2}(x^{\prime},x_{2},x_{3})\ \mathrm{d}x^{\prime}\Big),$
where $\bar{\bm{w}}_{h}$ is defined as the extension of ${\bm{w}}_{h}$ by 0 on $C_{R}\setminus\Omega$. From ${\bm{w}}_{h}\in C^{\infty}(\mathcal{M}_{h})$ we infer that ${\bm{A}}\in C^{\infty}(\mathcal{M}_{h})\cap H(\operatorname{{\bf curl}};\Omega)$, so that we may again apply the commuting projection operators (68). This yields
@@ 4.3 Discrete harmonic scalar potentials for $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$
The fields ${\bm{v}}_{h,i}:=\operatorname{{\bf grad}}\phi_{h,i}$, $i=1,\dots\beta_{2}$, form a basis of $\mathfrak{H}^{2}_{h}=\tilde{\mathfrak{H}}^{1}_{h,0}$. Moreover, they satisfy
@@ 4.4 Discrete harmonic vector potentials for $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$
$\left\{\begin{aligned} \langle\sigma_{h,i},\tau\rangle+\langle\operatorname{{\bf grad}}\tau,{\bm{A}}_{h,i}^{0}\rangle&=0,\ &&\quad\forall\ \tau\in\tilde{V}^{0}_{h,0},\\ \langle\operatorname{{\bf grad}}\sigma_{h,i},{\bm{v}}\rangle+\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{0},\operatorname{{\bf curl}}{\bm{v}}\rangle+\langle{\bm{p}}_{h,i},{\bm{v}}\rangle&=-\langle\operatorname{{\bf curl}}{\bm{A}}_{h,i}^{b},\operatorname{{\bf curl}}{\bm{v}}\rangle,\ &&\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0},\\ \langle{\bm{A}}_{h,i}^{0},{\bm{q}}\rangle&=0,\ &&\quad\forall\ {\bm{q}}\in\tilde{\mathfrak{H}}^{1}_{h,0}.\end{aligned}\right$. (84)
@@
We next observe that the flux functionals $\Phi_{i}$ from (40) are well-defined on $\tilde{\mathfrak{H}}^{2}_{h,0}$ since the latter is a subspace of $\tilde{V}^{2}_{0}=H_{0}(\operatorname{div};\Omega)$. We then have the following result.
@@
The fields ${\bm{w}}_{h,i}:=\operatorname{{\bf curl}}{\bm{A}}_{h,i}$, $i=1,\dots\beta_{1}$, form a basis of $\mathfrak{H}^{1}_{h}=\tilde{\mathfrak{H}}^{2}_{h,0}$. Moreover, their fluxes on the surfaces $\Sigma_{j}$ satisfy
$\Phi_{j}({\bm{w}}_{h,i})=\int_{\Sigma_{j}}{\bm{n}}\cdot{\bm{w}}_{h,i}=\delta_{i,j},\qquad 1\leq i,j\leq\beta_{1}.$ (85)
@@
From (82), it is clear that ${\bm{w}}_{h,i}=\operatorname{{\bf curl}}{\bm{A}}_{h,i}\in\tilde{V}^{2}_{h}$ satisfies $\operatorname{div}{\bm{w}}_{h,i}=0$. As for the boundary condition ${\bm{n}}\cdot{\bm{w}}_{h,i}=0$ on $\partial\Omega$, it is verified by writing
where we have used (83) and the commutation property (68) in the second equality. Here both term satisfies the boundary condition: the first one thanks to (46) and (66), and the second one as the curl of a field in $\tilde{V}^{1}_{0}=H_{0}(\operatorname{{\bf curl}};\Omega)$.
Thus, the fields ${\bm{w}}_{h,i}$ are in $\tilde{\mathfrak{Z}}^{2}_{h,0}$. To verify that they are in $\tilde{\mathfrak{H}}^{2}_{h,0}$ we must show that they also belong to $(\operatorname{{\bf curl}}\tilde{V}^{1}_{h,0})^{\perp}$, and for this we use the second equation from (84): taking first ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{h,i}$ shows that $\sigma_{h,i}=0$, then ${\bm{v}}={\bm{p}}_{h,i}$ yields ${\bm{p}}_{h,i}=0$, so that one finally obtains
$\langle{\bm{w}}_{h,i},\operatorname{{\bf curl}}{\bm{v}}\rangle=\langle\operatorname{{\bf curl}}({\bm{A}}^{0}_{h,i}+{\bm{A}}^{b}_{h,i}),\operatorname{{\bf curl}}{\bm{v}}\rangle=0,\quad\forall\ {\bm{v}}\in\tilde{V}^{1}_{h,0}$ (86)
which shows that ${\bm{w}}_{h,i}\in\tilde{\mathfrak{H}}^{2}_{h,0}$ indeed.
Finally, to evaluate the fluxes across a surface $\Sigma_{j}$ we use the fact that ${\bm{A}}^{0}_{h,i}\in H_{0}(\operatorname{{\bf curl}};\Omega)$, which yields $\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}^{0}_{h,i})=0$ according to Proposition 3.7, and (69) with ${\bm{a}}={\bm{A}}^{b}_{i}$: this gives
$\Phi_{j}({\bm{w}}_{h,i})=\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}^{b}_{h,i})=\Phi_{j}(\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{A}}^{b}_{i})=\int_{\Sigma_{j}}{\bm{n}}\cdot\operatorname{{\bf curl}}\tilde{\Pi}^{1}{\bm{A}}^{b}_{i}=\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}^{b}_{i}=\delta_{i,j}$
where the last step is again (46). This shows (85) and the linear independence of the fields ${\bm{w}}_{h,i}$. Their spanning properties follow from (72). ∎
Remark 4.4.
In the proof above, we further note that the first and last equations from (84) yield
@@
where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (70)–(71).
@@ 4.5 Solving a simpler discrete curl-curl problem
where we remind that $\tilde{\mathfrak{Z}}^{1}_{h,0}=\operatorname{{\bf grad}}\tilde{V}^{0}_{h,0}\oplus\tilde{\mathfrak{H}}^{1}_{h,0}$ is the kernel of the curl in $\tilde{V}^{1}_{h,0}$, see (70)–(71). Since the relations (90) characterize a unique field in $\tilde{V}^{1}_{h,0}$, this will show the existence and uniqueness of a solution. The fact that it coincides with the solution of (84) will follow from the fact that the latter satisfies the same relations, as seen in the proof of Theorem 4.2 and Remark 4.4.